What would the solubility of [Pb(OH)3]-(aq) be in 0.100 M Ba(OH)2 in the presence of excess Pb(OH)2(s)? Ksp Ba(OH)2 = 5.0 x 10-3 Ksp Pb(OH)2 = 1.43 x 10-20 Kf [Pb(OH)3]- = 3.8 x 1014
Answer #1

Ksp of Ba(OH)2 = 0.1 M

[OH-] = 2*0.1 = 0.2 M

Pb2+(aq) + 3OH- <----> [Pb(OH)3]-(aq); Kf = 3.8*1014 ---- Equilibrium 1

Pb(OH)2(s) <----> Pb2+(aq) + 2OH-(aq); Ksp = 1.43 x 10-20 ---- Equilibrium 2

Add the above two equilibria, you will get

Pb(OH)2(s) + OH- <----> [Pb(OH)3]-(aq); K = Ksp * Kf = 5.434*10-6

Now, K = [Pb(OH)3-]/[OH-]

i.e. 5.434*10-6 = [Pb(OH)3-]/0.2

i.e. [Pb(OH)3-] = 1.1*10-6 M